OpenStudy (anonymous):

Solve and round to three decimal places: e2x = 3/4 (

OpenStudy (anonymous):

To clarify e^(2x) = 3/4 ?

OpenStudy (anonymous):

if so, We want to isolate the x term. To so that, we gotta get rid of the e function Fortunately, we can undo the e by taking the natural log of it (ln) ln(e^(2x)) = ln(3/4) 2x = ln(3/4) x= ln(3/4) /2 = -0.144

OpenStudy (anonymous):

yes this is it, thank you